# atomic orbital API Application Programmer Interface

### v1

@@ 6. General case
$\Psi=\sum_{m=0}^{M}c_{m}\left|m\overline{m}\right\rangle$
Then the total energy takes the quadratic form
@@
$H_{km}=\left\langle k\overline{k}\left|\widehat{H}\right|m\overline{m}\right\rangle=2\delta_{km}(k|h|m)+(km|km)$
Introduce the Lagrangian
@@
with symmetric matrix coefficients $\varepsilon_{ij}$. It is a function of $\varphi_{0},\ldots,\varphi_{M},\varepsilon,c_{0},\ldots,c_{M}$ and $\varepsilon_{ij}$ with $i\leqslant j$. We calculate its gradient as follows. Firstly, the variational derivatives
$\frac{\delta\mathcal{L}}{\delta\varphi_{k}}$ $=c_{k}^{2}h\varphi_{k}+c_{k}\sum_{m=0}^{M}c_{m}J(km)\varphi_{m}-\sum_{m=0}^{M}\varepsilon_{km}\varphi_{m},\quad k=0,\ldots,M.$ (6.1)
$\frac{\partial\mathcal{L}}{\partial\varepsilon}$ $=-\frac{1}{4}\left(\sum_{m=0}^{M}c_{m}^{2}-1\right).$ (6.2)
$\frac{\partial\mathcal{L}}{\partial c_{k}}$ $=\frac{1}{2}\sum_{m=0}^{M}H_{km}c_{m}-\frac{1}{2}\varepsilon c_{k},\quad k=0,\ldots,M.$ (6.3)
$\frac{\partial\mathcal{L}}{\partial\varepsilon_{kk}}$ $=-\frac{1}{2}\left(\left\|\varphi_{k}\right\|^{2}-1\right),\quad k=0,\ldots,M,$ (6.4)
$\frac{\partial\mathcal{L}}{\partial\varepsilon_{ij}}$ $=-\int\varphi_{i}\varphi_{j},\quad 0\leqslant i<j\leqslant M.$ (6.5)
Now let us calculate the Hessian $d\nabla\mathcal{L}$.
$\sum_{j=0}^{M}\partial_{\varphi_{j}}\frac{\delta\mathcal{L}}{\delta\varphi_{k}}\delta\varphi_{j}=c_{k}^{2}h\delta\varphi_{k}-\sum_{m=0}^{M}\varepsilon_{km}\delta\varphi_{m}+c_{k}\sum_{m=0}^{M}c_{m}\left(J(\delta km)\varphi_{m}+J(k\delta m)\varphi_{m}+J(km)\delta\varphi_{m}\right),$ (6.6)
where $k=0,\ldots,M.$
$\frac{\partial}{\partial\varepsilon}\frac{\delta\mathcal{L}}{\delta\varphi_{k}}=0$
$\sum_{j=0}^{M}\frac{\partial}{\partial c_{j}}\frac{\delta\mathcal{L}}{\delta\varphi_{k}}\delta c_{j}=2c_{k}\delta c_{k}h\varphi_{k}+\delta c_{k}\sum_{m=0}^{M}c_{m}J(km)\varphi_{m}+c_{k}\sum_{m=0}^{M}\delta c_{m}J(km)\varphi_{m}$
@@
where we extended the orbital energy update by symmetry. Summing these equalities we obtain
$d\frac{\delta\mathcal{L}}{\delta\varphi_{k}}\delta w=c_{k}^{2}h\delta\varphi_{k}-\sum_{m=0}^{M}\varepsilon_{km}\delta\varphi_{m}+c_{k}\sum_{m=0}^{M}c_{m}\left(J(\delta km)\varphi_{m}+J(k\delta m)\varphi_{m}+J(km)\delta\varphi_{m}\right)\\ +2c_{k}\delta c_{k}h\varphi_{k}+\delta c_{k}\sum_{m=0}^{M}c_{m}J(km)\varphi_{m}+c_{k}\sum_{m=0}^{M}\delta c_{m}J(km)\varphi_{m}-\sum_{m=0}^{M}\delta\varepsilon_{km}\varphi_{m}$ (6.6)
$d\frac{\partial\mathcal{L}}{\partial\varepsilon}\delta w=-\frac{1}{2}\sum_{m=0}^{M}c_{m}\delta c_{m}$ (LHS2)
$\sum_{j=0}^{M}\partial_{\varphi_{j}}\frac{\partial\mathcal{L}}{\partial c_{k}}\delta\varphi_{j}=2c_{k}(\delta k|h|k)+\sum_{m=0}^{M}c_{m}(\delta km|km)+\sum_{m=0}^{M}c_{m}(k\delta m|km)$
@@
$\sum_{j=0}^{M}\frac{\partial}{\partial c_{j}}\frac{\partial\mathcal{L}}{\partial c_{k}}\delta c_{j}=\frac{1}{2}\sum_{m=0}^{M}H_{km}\delta c_{m}-\frac{1}{2}\varepsilon\delta c_{k}$
Summing these equalities we obtain
$d\frac{\partial\mathcal{L}}{\partial c_{k}}\delta w=2c_{k}(\delta k|h|k)+\sum_{m=0}^{M}c_{m}(\delta km|km)+\sum_{m=0}^{M}c_{m}(k\delta m|km)-\frac{1}{2}\delta\varepsilon c_{k}+\frac{1}{2}\sum_{m=0}^{M}H_{km}\delta c_{m}-\frac{1}{2}\varepsilon\delta c_{k}$ (LHS3)
Finally,
$d\frac{\partial\mathcal{L}}{\partial\varepsilon_{kk}}\delta w=-\int\varphi_{k}\delta\varphi_{k},\quad k=0,\ldots,M,$ (LHS4)
$d\frac{\partial\mathcal{L}}{\partial\varepsilon_{ij}}\delta w=-\int\delta\varphi_{i}\varphi_{j}-\int\varphi_{i}\delta\varphi_{j},\quad 0\leqslant i<j\leqslant M.$ (LHS5)
The right hand side is $-\nabla\mathcal{L}$. This finishes the detailed description of the Newton equation. Now we want to rewrite it in the self-consistent form, which will be suitable for an iterative procedure.
@@ 6.1. Self-consistent form
$\delta\varphi_{k}=\mathbb{F}_{k}(\delta\varphi_{0},\ldots,\delta\varphi_{M};w),\quad k=0,\ldots,M,$ (6.7)
where $w=(\varphi_{1},\ldots,\varepsilon_{M-1,M})$ is fixed at each Newton step, as follows.
@@
$\begin{pmatrix}0&\mathbf{c}^{T}\\ \mathbf{c}&\varepsilon-H\end{pmatrix}\begin{pmatrix}\delta\varepsilon\\ \delta\mathbf{c}\end{pmatrix}=\begin{pmatrix}-\frac{1}{2}\left(\sum_{m=0}^{M}c_{m}^{2}-1\right)\\ f\end{pmatrix}$
where: $\mathbf{c}=(c_{0},c_{1},\dots,c_{M})^{T}$ is a column vector, $\varepsilon-H$ is an $(M+1)\times(M+1)$ matrix. Here
$f_{k}=\sum_{m=0}^{M}H_{km}c_{m}-\varepsilon c_{k}+4c_{k}(\delta k|h|k)+2\sum_{m=0}^{M}c_{m}(\delta km|km)+2\sum_{m=0}^{M}c_{m}(k\delta m|km),\quad k=0,\ldots,M$
Introduce the matrix
$F_{kj}=\int\left(d\frac{\delta\mathcal{L}}{\delta\varphi_{k}}\delta w+\frac{\delta\mathcal{L}}{\delta\varphi_{k}}+\sum_{m=0}^{M}\delta\varepsilon_{km}\varphi_{m}+\sum_{m=0}^{M}\varepsilon_{km}\delta\varphi_{m}\right)\varphi_{j}=c_{k}^{2}(\delta k|h|j)\\ +c_{k}\sum_{m=0}^{M}c_{m}\left((\delta km|jm)+(k\delta m|jm)+(j\delta m|km)\right)+2c_{k}\delta c_{k}(k|h|j)+\sum_{m=0}^{M}(c_{k}\delta c_{m}+\delta c_{k}c_{m})(jm|km)\\ +c_{k}^{2}(k|h|j)+c_{k}\sum_{m=0}^{M}c_{m}(jm|km)-\underbrace{\sum_{m=0}^{M}\varepsilon_{km}\int\varphi_{j}\varphi_{m}}_{=\varepsilon_{jk}}$
@@
where $\mathcal{E}=(\varepsilon_{kj})$, $X=(\delta\varepsilon_{kj})$ is symmetric and $Y=\left(\int\delta\varphi_{j}\varphi_{k}\right)$ is antisymmetric. As long as the spectra of $\mathcal{E}$ and $-\mathcal{E}$ are disjoint, there exist a unique solution $X$. For this it is enough to have all the eigenvalues of $\mathcal{E}$ being negative. One naturally anticipates the orbital energies to be negative. Taking the transpose of both sides and accounting for $X^{T}=X$, $Y^{T}=-Y$ one obtains
@@
The equation for $Y$ is a Sylvester equation, which can be solved by the Bartels–Stewart algorithm implemented in many software packages. Its computational cost is $\mathcal{O}\left(M^{3}\right)$ arithmetical operations, that can be viewed negligible. Once $Y$ is found, we solve for $X$:
@@
It is left to precondition the first $M+1$ equations in the Newton system
@@
as follows
@@
Finally,
$\delta\varphi_{k}=-\left(1+\frac{2\delta c_{k}}{c_{k}}\right)\varphi_{k}-R_{k}\mathfrak{F}_{k}$ (6.8)
@@
and the convolution operator $R_{k}$ is defined as above. This finishes the description of $\mathbb{F}_{k}$ in (6.7).
@@ 6.2. Numerics
Figure 2. Three determinants ground state of the hydrogen molecule with the coefficients $c_{0}=0.99253,c_{1}=-0.10718$ and $c_{2}=-0.05829$, accordingly. The corresponding energy is $-1.15949$.
The default calculation settings: maximum 15 iterations for the inner loop related to the treatment of each Newton’s equation. We use DIIS with maximum 3 previous iterations. The initial trust radius is set to 1.0 and modified by dividing in two every time the new iteration results in a bigger energy value. In addition, we modify the diagonal values of the orbital energies by multiplying by 10. We initialize the orbital energy matrix by $\mathds{1}$ at the beginning. One may reset the orbital energy matrix back to $-\mathds{1}$ at the end of the first few iterations, in order to avoid unwanted saddle points.
[…]
Figure 3. Newton energy optimization step.
After each iteration the state is normalized using Löwdin transform. In addition, CI coefficients are optimized after the orthonormalization, see Figure 3. These two sub-steps are very cheap comparing to solving Newton’s system. The converged result for the two orbital MCSCF for the helium atom with the multiwavelet threshold $\varepsilon_{\text{MRA}}=10^{-5}$ is $c_{0}=0.99793,c_{1}=-0.06430$ and orbitals symmetric with respect to origin, shown in Figure 1. The corresponding total energy is $-2.87799$. It is worth to compare with [16], where the authors report a ground state energy for Helium of -2.90372438136211 a.u. using their free iterative complement interaction (ICI) method. This result agrees with other high-precision calculations [3]. It’s worth noting that while this is a theoretical calculation, it’s considered extremely accurate and is often used as a benchmark for experimental measurements and other theoretical methods.
For H2 molecule we refer to [14]: the equilibrium internuclear distance $R_{\text{nuc}}=1.4010784$, the nuclear repulsion corrected total energy $E_{\text{el}}+1/R_{\text{nuc}}$ is -1.17447. These values are treated as the reference in the current work. In [17] the total energy of $-1.174475931399$ hartree at equilibrium $R_{\text{nuc}}=1.4011$ is reported. Our 3 configurations result is the energy $-1.15949$ with the coefficients $c_{0}=0.99253,c_{1}=-0.10718$ and $c_{2}=-0.05829$. The corresponding MCSCF orbitals are shown in Figure 2.
The MCSCF approximation approaches the corresponding exact wave function very slowly with the increase of determinants, see Figures 4, 5. For comparison we also included the calculations with B3LYP functional. The DFT method provides a lower value than $E_{\text{exact}}$ for He, and a higher value than $E_{\text{exact}}$ for H2.
@@
It is worth to notice, that we didn’t rely on any chemical intuition in order to choose proper initial guesses for the two quantum systems we regarded. Instead, we first ran single electron calculations starting with orbitals combining randomly localized gaussian functions. The converged ionic simulations were later used for initialization of the Newton treatment of MCSCF. We follow the same initial guess strategy for the excited states, considered below. Alternatively, one can use spherical harmonics appearing in hydrogen-like atoms. However, it is not clear how this initialization approach could be extended to bigger molecules. Therefore, we stick to the randomized initial calculations.

### v2

@@ 5. General case
(5.1) $\Psi=\sum_{m=0}^{M}c_{m}\left|m\overline{m}\right\rangle$
with orthonormal spatial orbitals $\varphi_{m}$ and opposite spins. This is a natural expansion in the sense that it does not lead to loss of generality, as was first explained by Löwdin and Shull [13]. In other words, a wave function $\Psi$ of rank $M+1$ for a closed shell molecule can always be rewritten in the form (5.1). In Appendix B we provide a full proof of this claim from the first principles, namely, from the closed-shell assumption $\hat{S}^{2}\Psi=0.$ It is worth to point out that the representation (5.1) was derived by Löwdin and Shull, but not from the wave function zero spin condition. They assumed that the wave function already has a representation, where each Slater determinant has two spin orbitals with opposite spins. Then they deduced the orthogonality of the corresponding spatial orbitals. Therefore, we complement their derivation significantly.
For the natural expansion (5.1) it is easy to check that the total energy takes the quadratic form
@@
$H_{km}=\left\langle k\overline{k}\left|\widehat{H}\right|m\overline{m}\right\rangle=2\delta_{km}(k|h|m)+(km|km).$
We introduce the Lagrangian
@@
with symmetric matrix coefficients $\varepsilon_{ij}$. It is a function of the variables $\varphi_{0},\ldots,\varphi_{M},\varepsilon,c_{0},\ldots,c_{M}$ and $\varepsilon_{ij}$ with $i\leqslant j$. We compute its gradient $\nabla\mathcal{L}$ and organize its components as follows. First, the variational derivatives and then the partial derivatives:
(5.2) $\frac{\delta\mathcal{L}}{\delta\varphi_{k}}$ $=c_{k}^{2}h\varphi_{k}+c_{k}\sum_{m=0}^{M}c_{m}J(km)\varphi_{m}-\sum_{m=0}^{M}\varepsilon_{km}\varphi_{m},\quad k=0,\ldots,M,$
(5.3) $\frac{\partial\mathcal{L}}{\partial\varepsilon}$ $=-\frac{1}{4}\left(\sum_{m=0}^{M}c_{m}^{2}-1\right),$
(5.4) $\frac{\partial\mathcal{L}}{\partial c_{k}}$ $=\frac{1}{2}\sum_{m=0}^{M}H_{km}c_{m}-\frac{1}{2}\varepsilon c_{k},\quad k=0,\ldots,M,$
(5.5) $\frac{\partial\mathcal{L}}{\partial\varepsilon_{kk}}$ $=-\frac{1}{2}\left(\left\|\varphi_{k}\right\|^{2}-1\right),\quad k=0,\ldots,M,$
(5.6) $\frac{\partial\mathcal{L}}{\partial\varepsilon_{ij}}$ $=-\int\varphi_{i}\varphi_{j},\quad 0\leqslant i<j\leqslant M.$
Now let us calculate the Hessian $d\nabla\mathcal{L}$. We differentiate (5.2) with respect to orbitals as
(5.7) $\sum_{j=0}^{M}\partial_{\varphi_{j}}\frac{\delta\mathcal{L}}{\delta\varphi_{k}}\delta\varphi_{j}=c_{k}^{2}h\delta\varphi_{k}-\sum_{m=0}^{M}\varepsilon_{km}\delta\varphi_{m}+c_{k}\sum_{m=0}^{M}c_{m}\left(J(\delta km)\varphi_{m}+J(k\delta m)\varphi_{m}+J(km)\delta\varphi_{m}\right),$
where $k=0,\ldots,M.$ Next we continue differentiating (5.2) as
$\frac{\partial}{\partial\varepsilon}\frac{\delta\mathcal{L}}{\delta\varphi_{k}}=0,$
$\sum_{j=0}^{M}\frac{\partial}{\partial c_{j}}\frac{\delta\mathcal{L}}{\delta\varphi_{k}}\delta c_{j}=2c_{k}\delta c_{k}h\varphi_{k}+\delta c_{k}\sum_{m=0}^{M}c_{m}J(km)\varphi_{m}+c_{k}\sum_{m=0}^{M}\delta c_{m}J(km)\varphi_{m},$
@@
where we extend the orbital energy update by symmetry. Summing these equalities, we obtain
(5.7) $d\frac{\delta\mathcal{L}}{\delta\varphi_{k}}\delta w=c_{k}^{2}h\delta\varphi_{k}-\sum_{m=0}^{M}\varepsilon_{km}\delta\varphi_{m}+c_{k}\sum_{m=0}^{M}c_{m}\left(J(\delta km)\varphi_{m}+J(k\delta m)\varphi_{m}+J(km)\delta\varphi_{m}\right)\\ +2c_{k}\delta c_{k}h\varphi_{k}+\delta c_{k}\sum_{m=0}^{M}c_{m}J(km)\varphi_{m}+c_{k}\sum_{m=0}^{M}\delta c_{m}J(km)\varphi_{m}-\sum_{m=0}^{M}\delta\varepsilon_{km}\varphi_{m},$
which are the first $M+1$ lines of the left had side of the Newton system. Next, we have
(LHS2) $d\frac{\partial\mathcal{L}}{\partial\varepsilon}\delta w=-\frac{1}{2}\sum_{m=0}^{M}c_{m}\delta c_{m}.$
We compute
$\sum_{j=0}^{M}\partial_{\varphi_{j}}\frac{\partial\mathcal{L}}{\partial c_{k}}\delta\varphi_{j}=2c_{k}(\delta k|h|k)+\sum_{m=0}^{M}c_{m}(\delta km|km)+\sum_{m=0}^{M}c_{m}(k\delta m|km),$
@@
and
$\sum_{j=0}^{M}\frac{\partial}{\partial c_{j}}\frac{\partial\mathcal{L}}{\partial c_{k}}\delta c_{j}=\frac{1}{2}\sum_{m=0}^{M}H_{km}\delta c_{m}-\frac{1}{2}\varepsilon\delta c_{k}.$
Summing these equalities, we obtain
(LHS3) $d\frac{\partial\mathcal{L}}{\partial c_{k}}\delta w=2c_{k}(\delta k|h|k)+\sum_{m=0}^{M}c_{m}(\delta km|km)+\sum_{m=0}^{M}c_{m}(k\delta m|km)-\frac{1}{2}\delta\varepsilon c_{k}+\frac{1}{2}\sum_{m=0}^{M}H_{km}\delta c_{m}-\frac{1}{2}\varepsilon\delta c_{k}.$
Finally, the last parts of the Hessian have the form
(LHS4) $d\frac{\partial\mathcal{L}}{\partial\varepsilon_{kk}}\delta w=-\int\varphi_{k}\delta\varphi_{k},\quad k=0,\ldots,M,$
(LHS5) $d\frac{\partial\mathcal{L}}{\partial\varepsilon_{ij}}\delta w=-\int\delta\varphi_{i}\varphi_{j}-\int\varphi_{i}\delta\varphi_{j},\quad 0\leqslant i<j\leqslant M.$
Expressions (5.7)-(LHS5) define the Hessian and form the left-hand side of the Newton equation. The right-hand side is -$\nabla\mathcal{L}$ with the gradient components given by (5.2)-(5.6). This completes the detailed description of the Newton equation (2.3) with $\mathcal{R}=\nabla\mathcal{L}$. We now rewrite it in the self-consistent form (2.4), which is suitable for an iterative procedure.
@@ 5.1. Self-consistent form
(5.8) $\delta\varphi_{k}=\mathbb{F}_{k}(\delta\varphi_{0},\ldots,\delta\varphi_{M};w),\quad k=0,\ldots,M,$
where $w=(\varphi_{1},\ldots,\varepsilon_{M-1,M})$ is held fixed at each Newton step, as follows. We can isolate a subsystem of the form
@@
in Newton’s equation. Shortly,
$\begin{pmatrix}0&\mathbf{c}^{T}\\ \mathbf{c}&\varepsilon-H\end{pmatrix}\begin{pmatrix}\delta\varepsilon\\ \delta\mathbf{c}\end{pmatrix}=\begin{pmatrix}-\frac{1}{2}\left(\sum_{m=0}^{M}c_{m}^{2}-1\right)\\ f\end{pmatrix}$,
where $\mathbf{c}=(c_{0},c_{1},\dots,c_{M})^{T}$ is a column vector and $\varepsilon-H$ is an $(M+1)\times(M+1)$ matrix. On the right-hand side we have a column vector $f$ with the components
$f_{k}=\sum_{m=0}^{M}H_{km}c_{m}-\varepsilon c_{k}+4c_{k}(\delta k|h|k)+2\sum_{m=0}^{M}c_{m}(\delta km|km)+2\sum_{m=0}^{M}c_{m}(k\delta m|km),\quad k=0,\ldots,M.$
Further on, we introduce the matrix
$F_{kj}=\int\left(d\frac{\delta\mathcal{L}}{\delta\varphi_{k}}\delta w+\frac{\delta\mathcal{L}}{\delta\varphi_{k}}+\sum_{m=0}^{M}\delta\varepsilon_{km}\varphi_{m}+\sum_{m=0}^{M}\varepsilon_{km}\delta\varphi_{m}\right)\varphi_{j}=c_{k}^{2}(\delta k|h|j)\\ +c_{k}\sum_{m=0}^{M}c_{m}\left((\delta km|jm)+(k\delta m|jm)+(j\delta m|km)\right)+2c_{k}\delta c_{k}(k|h|j)+\sum_{m=0}^{M}(c_{k}\delta c_{m}+\delta c_{k}c_{m})(jm|km)\\ +c_{k}^{2}(k|h|j)+c_{k}\sum_{m=0}^{M}c_{m}(jm|km)-\underbrace{\sum_{m=0}^{M}\varepsilon_{km}\int\varphi_{j}\varphi_{m}}_{=\varepsilon_{jk}}.$
@@
where $\mathcal{E}=(\varepsilon_{kj})$, $X=(\delta\varepsilon_{kj})$ is symmetric and $Y=\left(\int\delta\varphi_{j}\varphi_{k}\right)$ is antisymmetric. As long as the spectra of $\mathcal{E}$ and -$\mathcal{E}$ are disjoint, there exists a unique solution $X$. For this it is enough to have all the eigenvalues of $\mathcal{E}$ being negative. One naturally anticipates the orbital energies to be negative. Taking the transpose of both sides and accounting for $X^{T}=X$, $Y^{T}=-Y$ one obtains
@@
The equation for $Y$ is a Sylvester equation, which can be solved by the Bartels–Stewart algorithm implemented in many software packages. Its computational cost is $\mathcal{O}\left(M^{3}\right)$ arithmetical operations, which can be viewed as negligible. Once $Y$ has been found, we solve for $X$:
@@
It remains to precondition the first $M+1$ equations in the Newton system
@@
by first isolating the term $\frac{c_{k}^{2}}{2}\Delta-\varepsilon_{kk}$ in
@@
and then applying the resolvent $R_{k}$. This leads to
(5.9) $\delta\varphi_{k}=-\left(1+\frac{2\delta c_{k}}{c_{k}}\right)\varphi_{k}-R_{k}\mathfrak{F}_{k},$
@@
and the convolution operator $R_{k}$ is defined in (2.1). This finishes the description of $\mathbb{F}_{k}$ in (5.8).
@@ 5.2. Numerics
Figure 3. Three-determinant ground state of the hydrogen molecule with the coefficients $c_{0}=0.99253,c_{1}=-0.10718$ and $c_{2}=-0.05829$, accordingly. The corresponding energy is 1.15949-1.15949.
The default calculation settings are as follows: a maximum of 15 iterations for the inner loop associated with the solution of each Newton equation. We use DIIS with maximum 3 previous iterations. The initial trust radius is set to 1.0 and modified by dividing by two every time the new iteration results in a higher energy value. In addition, we modify the diagonal values of the orbital energies by multiplying by 10. We initialize the orbital energy matrix by $\mathds{1}$ at the beginning. One may reset the orbital energy matrix back to $\mathds{1}$ at the end of the first few iterations, in order to avoid unwanted saddle points.
After each iteration the state is normalized using Löwdin transform. In addition, CI coefficients are optimized after the orthonormalization, see Figure 1. These two sub-steps are very cheap compared to solving the Newton system. The converged result for the two orbital MCSCF for the helium atom with the multiwavelet threshold $\varepsilon_{\text{MRA}}=10^{-5}$ is $c_{0}=0.99793,c_{1}=-0.06430$ and orbitals symmetric with respect to origin, shown in Figure 2. The corresponding total energy is 2.87799-2.87799. It is worth comparing with [15], where the authors report a ground state energy for Helium of -2.90372438136211 a.u. using their free iterative complement interaction (ICI) method. This result agrees with other high-precision calculations [1]. It is worth noting that, while this is a theoretical calculation, it is considered extremely accurate and is often used as a benchmark for experimental measurements and other theoretical methods.
For H2 molecule, we refer to [12]: the equilibrium internuclear distance $R_{\text{nuc}}=1.4010784$, the nuclear repulsion corrected total energy $E_{\text{el}}+1/R_{\text{nuc}}$ is -1.17447. These values are used as reference values in the present work. In [20] the total energy of 1.174475931399-1.174475931399 hartree at equilibrium $R_{\text{nuc}}=1.4011$ is reported. Our three-configuration result yields the energy 1.15949-1.15949 with the coefficients $c_{0}=0.99253,c_{1}=-0.10718$ and $c_{2}=-0.05829$. The corresponding MCSCF orbitals are shown in Figure 3.
The MCSCF approximation approaches the corresponding exact wave function very slowly as the number of determinants increases, see Figures 4, 5. For comparison, we also include the calculations with B3LYP functional. The DFT method provides a lower value than $E_{\text{exact}}$ for He, and a higher value than $E_{\text{exact}}$ for H2.
@@
It is worth noting that we did not rely on any chemical intuition in order to choose proper initial guesses for the two quantum systems we regarded. Instead, we first ran single electron calculations starting with orbitals combining randomly localized Gaussian functions. The converged ionic simulations were later used to initialize the Newton treatment of MCSCF. We follow the same initial guess strategy for the excited states considered below. Alternatively, one can use spherical harmonics appearing in hydrogen-like atoms. However, it is not clear how this initialization approach can be extended to larger molecules. Therefore, we adopt randomized initial calculations.
5.3. Single-determinant problem
For the Hartree–Fock problem with $M=0$, if the orbital $\varphi_{0}$ is normalized after each Newton step, then
$\int\varphi_{0}\delta\varphi_{0}=0,$
which simplifies the scalar subsystem
$\delta\varepsilon=(\delta 0|h|0)+3(0\delta 0|00)+(0|h|0)+(00|00)-\varepsilon.$
Moreover, one can reset the Lagrange multiplier before running Newton’s inner loop as
$\varepsilon=(0|h|0)+(00|00),$
in agreement with the Hartree–Fock equation. Under these assumptions, the Newton system reduces to
[…]
$\delta\varphi$ $=-\varphi+\left(-\frac{\Delta}{2}-\varepsilon\right)^{-1}\Big[\delta\varepsilon\varphi-\left(V_{\text{nuc}}+J(00)\right)\left(\delta\varphi+\varphi\right)-2J(0\delta 0)\varphi\Big].$
Furthermore, if one approximates the solution of the Newton system by a single inner iteration,
$\delta\varphi=0\;\;\mapsto\;\;\delta\varepsilon=0\;\;\mapsto\;\;\delta\varphi,$
then the update reduces to
$\varphi+\delta\varphi=-\left(-\frac{\Delta}{2}-\varepsilon\right)^{-1}\Big[\left(V_{\text{nuc}}+J(00)\right)\varphi\Big],$
which coincides with a well-established Green’s function iteration scheme [7, 9]. This observation highlights the fundamental role of the resolvent operator and provides further insight into the efficiency of multiwavelet-based methods in electronic structure calculations.